OpenStudy (isaiahriveraeduardo):

A department store's customer parking lot has 4 rows with an equal number of parking spots in each row. The lot also has 15 parking spots for store employees. If c cars can be parked in each of the 4 main rows of the parking lot, what is the expression for the maximum number of cars that can be parked in the parking lot?

OpenStudy (isaiahriveraeduardo):

...

OpenStudy (queenie):

would this happen to be multiple choice

OpenStudy (isaiahriveraeduardo):

yes

OpenStudy (queenie):

answer choices please

OpenStudy (isaiahriveraeduardo):

15c − 4 c(15 − 4) 4c + 15 4(c + 15)

OpenStudy (queenie):

I believe it is 4c + 15

OpenStudy (welshfella):

@queenie please explain how you got that answer.

OpenStudy (welshfella):

direct answers are not allowed

OpenStudy (queenie):

there are 4 rows with equal number of cars.....+ the 15 for the employees. Lets say that each row has 10 cars.... 4(10) + 15 = 40 + 15 = 55 cars understand ?